OpenStudy (anonymous):

at 1 degrees Celcius. under a pressure of 20kPa a given sample of gas occupies 75.0 liters. The gas is compressed to a final volume of 30.0 liters at 0 degrees Celcius .What is the final pressure in kPa?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (cuanchi):

P1V1/T1=P2V2/T2 P2= P1V1T2/V2T1 convert the temperatures from C to Kelvin and solve for P2 K=C+273

OpenStudy (anonymous):

@Cuanchi

OpenStudy (anonymous):

@Cuanchi

OpenStudy (anonymous):

@Cuanchi

OpenStudy (anonymous):

@perl